# NeurIPS 2025 Workshop on Time Series for Health

### v1

@@ 4 Experiments
We apply Fitted Q-Evaluation (FQE) off-policy evaluation (OPE) zhang2022off ; uehara2022review for all models in Table 3. For each model, we compute test set Q-values and fit a Gaussian to summarize their distribution. We compare model optimal policy with the clinician actions. The results show that block discrete models perform as well as continuous variable for $vp_{2}$. Stepwise directional with a larger max_step, has the largest 25.1% survivial improvement probability. Moreover, incorporating an LSTM with recurrent experience reply yields a survival improvement probability of 17.4%. This shows the benefits of capturing cross-time dependencies.
Table 3: FQE off-policy evaluation with the best $\alpha$ from validation. We define $\Delta$ Surv. Imp. Prob. (%) (Survival Improvement Probability) as the percentage of model recommended policy reaching higher reward value above the clinician mean reward minus 50% (PP(Model > Cli. Mean) - $50\%$), and $\Delta$ Q Mean (Mean Improvement) as the shift in the mean of fitted Gaussian for Model vs Clinician.
Model $\alpha$ Config $\Delta$ Surv. Imp. Prob. (%) $\Delta$ Q Mean Model M. Cli. M. Cohen’s d
@@
5 Conclusion and Future Work
This paper establishes action-space design as a critical bridge between reinforcement learning optimization and clinical implementation in ICU settings. By developing a novel action space, we convert theoretical optimal policies into actionable clinical protocols. Our proposed dual vasopressor control system for septic shock demonstrates that recurrent experience replay effectively captures the complex temporal dependencies inherent in ICU patient responses, achieving superior performance of over 15% survival improvement probability. These contributions move towards a blueprint for CDSS systems that clinicians could interpret and trust, and towards AI systems that not only compute optimal actions, but deliver them in forms that integrate into critical medical workflows.
Appendix A Fitted Q-Evaluation Illustrations
@@
Appendix B Detailed Experiment Results
@@
Dual Blk. Disc. 3 bins 0.000 71.4 1.734 0.123 55.2 57.1
Dual Blk. Disc. 5 bins 0.000 96.3 2.383 0.265 39.7 26.4
Dual Blk. Disc. 10 bins 0.000 92.3 4.673 0.309 42.4 13.3
@@
Dual Blk. Disc. 3 bins 0.001 82.2 1.614 0.110 46.9 55.5
Dual Blk. Disc. 5 bins 0.001 93.6 2.281 0.217 41.3 27.2
Dual Blk. Disc. 10 bins 0.001 96.6 4.445 0.292 39.1 14.1
@@
Dual Blk. Disc. 3 bins 0.010 76.8 1.576 0.127 53.1 56.3
Dual Blk. Disc. 5 bins 0.010 84.9 2.091 0.119 43.9 40.3
Dual Blk. Disc. 10 bins 0.010 92.8 4.072 0.245 41.4 16.6
@@
Appendix C Model Implementation and Training
C.1 Double Q-learning
@@
C.2 Binary vp1 Model
The Binary $vp_{1}$ model represents the simplest action space, treating vasopressin administration as a binary decision (on/off). The vp2 levels are used as part of the state. Training employs standard Q-learning TD loss with conservative logsumexp regularization over both action choices to prevent overestimation of out-of-distribution actions. The model processes states through fully connected networks with ReLU activations, using double Q-learning with target networks for stability.
C.3 Dual Mixed Model
@@
C.4 Dual Block Discrete Model
@@
C.5 Dual Stepwise Model
@@
C.6 LSTM Block Discrete Model
@@
Appendix D Model Architectures
@@
Appendix E Hyperparameter Configuration
E.1 Shared Hyperparameters (All Models)
@@
Epochs 100 Training epochs
Validation Batches 10 Batches for validation evaluation
Random Seed 42 For reproducibility
E.2 Model-Specific Configurations
@@
Appendix F Offline Policy Evaluation with Fitted Q-Evaluation (FQE)
To evaluate the learned policies against historical clinician decisions without online deployment, we employed Fitted Q-Evaluation (FQE) with Gaussian distribution modeling. This approach computes Q-values for both the learned policy and clinician actions on the held-out test set, enabling direct comparison of expected returns. For each trained model, we extract Q-values by evaluating state-action pairs from the test trajectories where actions are either from the learned policy (model Q-values) or from historical clinician decisions (clinician Q-values). The resulting Q-value distributions are then fitted with Gaussian distributions using maximum likelihood estimation, providing parametric representations characterized by mean ($\mu$) and standard deviation ($\sigma$) for both the model and clinician policies. This parametric approach enables computation of key metrics including the mean improvement gap (difference in expected returns), probability of improvement (likelihood that model Q-values exceed clinician Q-values), and effect size (Cohen’s d) to quantify the magnitude of policy differences.
The FQE analysis also generates visualizations including overlapping histograms with fitted Gaussian curves, cumulative distribution functions (CDFs) showing the probability mass distribution of Q-values, and improvement probability curves that illustrate the likelihood of the model outperforming various Q-value thresholds. The probability of improvement at the clinician’s mean Q-value serves as a critical metric, indicating how often the learned policy is expected to achieve better outcomes than historical treatment decisions. Additionally, the analysis computes Cohen’s d as a standardized effect size measure. The larger the value of Cohen’s d, the more meaningful are the difference between the means of two groups. This framework provides quantitative evidence for policy improvement while accounting for the inherent uncertainty in Q-value estimates. The FQE analysis was performed for policy performance across different model architectures and hyperparameter configurations. The FQE numerical results are given in Table 3. The histogram and sigmoid CDF visualizations are shown in Appendix A.

### v2

@@ 4.1 Learnability of Clinically Aligned Action Spaces
Fitted Q-Evaluation  [^7] See Appendix B and C for detailed FQE descriptions and figures .. FQE is applied as an off-policy evaluation (OPE) zhang2022off ; uehara2022review method for all models. For each model, we compute test-set Q-values and fit a Gaussian to summarize their distribution. We compare model optimal policy with the clinician actions. As shown in Table 3, block discrete models perform as well as continuous variable for $vp_{2}$. Stepwise directional with a larger max_step has the largest $25.1\%$ probability of improved rewards. Moreover, incorporating an LSTM with recurrent experience reply yields a $17.4\%$ probability of improved rewards, which shows the benefits of capturing cross-time dependencies.
Table 3: FQE OPE Results. PIR (Probability of Improved Rewards) is defined as the probability of model recommended action reaching a higher reward value than the clinician mean reward, minus $50\%$ (PP(Model > Cli. Mean) - $50\%$); $\Delta$ Q Mean is the shift in the mean of fitted Gaussian for Model vs Clinician.
Model $\alpha$ Config PIR $\Delta$ Q Mean Mod. Mean Cli. Mean Cohen’s d
@@
4.2 Performance with Full Clinical Reward and Weighted Importance Sampling (WIS)
To show that the action space design concretely improves model performance, we implement the full reward function in kalimouttou2025optimal , and leverage WIS for Offline-Policy Evaluation (OPE).
Comprehensive Reward Function. Following kalimouttou2025optimal , we implement a comprehensive reward function with adjusted rewards that indicate patient progress. The reward is composed of a base survival reward, a set of benefit for improved clinical parameters (such as improved mean blood pressure or lactate levels), reward for decreased norepinephrine usage, and mortality penalty  [^8] See Appendix D for more detailed reward description..
Weighted Importance Sampling. OPE is performed on the test-set trajectories using. First, we identify optimal actions recommended by our model and train a softmax model on both the target policy (optimal model actions, $\pi_{m}$) and baseline policy (clinician actions, $\pi_{c}$). Then, the trajectory-level WIS coefficients are computed as specified in the first line of Equation 2 to estimate the trajectory-level expected reward $R^{(i)}_{\text{traj}}$. Lastly, the overall WIS expected reward is computed using the renormalization shown on the second line of Equation 2.
$R^{(i)}_{\text{traj}}$ $=T\cdot\frac{\sum_{t=0}^{T}w_{t}R_{t}}{\sum_{t=0}^{T}w_{t}},$ $\text{where }w_{t}=\frac{\pi_{m}(a_{t}^{\text{cli.}}\mid s_{t})}{\pi_{c}(a_{t}^{\text{cli.}}\mid s_{t})},$ (2)
$R_{\text{WIS}}$ $=\frac{\sum_{i=0}^{N}\tilde{w}_{i}R_{\text{traj}}^{(i)}}{\sum_{i=0}^{N}\tilde{w}_{i}},$ $\text{where }\tilde{w}_{i}=\frac{1}{T}\sum_{t=0}^{T}w_{t}.$
We also calculate the trajectory-level WIS by computing the trajectory-level weight as the product of transition level weights across the trajectory time-steps. This alternative evaluation resulted in similar ranges of WIS improvements as in kalimouttou2025optimal . For our clinically-aligned discrete action spaces, we found the two-step WIS method described in Equation 2 to be reliable and stable.
Model implementation, architecture, and hyperparameters are the same as Section 4.1 except that we remove conservative Q-learning ($\alpha=0.0$), and train all models for $500$ epochs.
Results. As shown in Table 4, finer discretization in the action space yields better performance, and adding recurrent LSTMs with experience replay results in further substantial improvements. Both the Dual BD and LSTM BD models show over $3\times$ improved WIS expected rewards, compared with the the Dual Mixed model where $vp_{2}$ is a continuous variable.
Table 4: Importance Sampling (IS) and Weighted Importance Sampling (WIS) OPE Results. We show transition-level (Trans. Lvl) average reward $E(R)$, IS, WIS, and corrsponding trajectory-level (Traj. Lvl) average reward. We also present the improvements in trajectory-level WIS (Imp. WIS) between model recommended actions and clinician actions. Lastly, we show the $95\%$ Confidence Interval (CI) for WIS improvement obtained by bootstrapping across test-set trajectories.
[…]
$E[R]$ IS WIS $E[R]$ IS WIS (Mod.-Cli.)
Binary $vp_{1}$ 2.13 2.10 2.15 112.01 110.51 117.54 5.53 - [2.0, 9.1]
[…]
5 Conclusion
[…]
Appendix
Appendix A Data Source and Processing
The patient features extracted from the eICU and MIMIC-IV databases serve as the state representation for the reinforcement learning (RL) algorithm, and the action space consists of vasopressin administration (binary) and norepinephrine dosing in the continuous range $(0,0.5]$ (mcg/kg/min). The data are preprocessed such that missing values are imputed using forward filling, and norepinephrine dosages are clipped to the valid clinical range $(0,0.5]$. We include unique patients experiencing their first episode of septic shock, defined using the Sepsis-3 criteria [27] and who were already receiving norepinephrine at the moment shock onset was identified. Shock onset could occur either while the patient was still in the emergency department or after ICU admission.
Appendix B Fitted Q-Evaluation (FQE)
To evaluate the learned policies against historical clinician decisions without online deployment, we employ Fitted Q-Evaluation (FQE) with Gaussian distribution modeling. This approach computes Q-values for both the learned policy and clinician actions on the held-out test set, enabling direct comparison of expected returns. For each trained model, we extract Q-values by evaluating state-action pairs from the test trajectories where actions are either from the learned policy (model Q-values) or from historical clinician decisions (clinician Q-values). The resulting Q-value distributions are then fitted with Gaussian distributions using maximum likelihood estimation, providing parametric representations characterized by mean ($\mu$) and standard deviation ($\sigma$) for both the model and clinician policies. This parametric approach enables computation of key metrics including the mean improvement gap (difference in expected returns), probability of improvement (likelihood that model Q-values exceed clinician Q-values), and effect size (Cohen’s d) to quantify the magnitude of policy differences.
The FQE analysis also generates visualizations including overlapping histograms with fitted Gaussian curves, cumulative distribution functions (CDFs) showing the probability mass distribution of Q-values, and improvement probability curves that illustrate the likelihood of the model outperforming various Q-value thresholds. The probability of improvement at the clinician’s mean Q-value serves as a critical metric, indicating how often the learned policy is expected to achieve better outcomes than historical treatment decisions. Additionally, the analysis computes Cohen’s d as a standardized effect size measure. The larger the value of Cohen’s d, the more meaningful are the difference between the means of two groups. This framework provides quantitative evidence for policy improvement while accounting for the inherent uncertainty in Q-value estimates. The FQE analysis was performed for policy performance across different model architectures and hyperparameter configurations. The FQE numerical results are given in Table 3. The histogram and sigmoid CDF visualizations are shown in Appendix C.
Appendix C Fitted Q-Evaluation Illustrations
@@
Appendix D Comprehensive Reward Function
We directly follow the OVISS [7] study to implement a full and comprehensive reward function with adjusted rewards. The reward measures patient progress and response to vasopressor administration in a quantifiable manner. The reward definition involves calculating an "adjusted reward" for each patient at different time points, which is a composite measure reflecting both survival and physiological benefits.
  • Survival Reward. Each patient receives a base reward of +1+1 at every time step, reflecting the fundamental importance of maintaining survival.
  • Death Penalty. If in-hospital mortality occurs, a penalty of 20-20 is applied at the terminal state of the trajectory. This large negative reward emphasizes the critical nature of mortality when evaluating treatment effectiveness.
  • Clinical Improvement Rewards. Additional positive rewards are assigned when key physiological markers improve within prespecified future windows:
    • Lactate Decrease (+1): A decrease in lactate levels within the next six hours, indicating improved metabolic status.
    • Mean Blood Pressure Increase (+1): MBP rising to $\geq 65$ mmHg within the next four hours, for cases where the initial MBP is below this threshold, reflecting improved hemodynamic stability.
    • SOFA Score Decrease (+3): A reduction in SOFA score within the next six hours, highlighting improved multi-organ function.
    • Reduced Norepinephrine Usage (+1): A decrease in norepinephrine dosage within the next four hours, suggesting better cardiovascular stability.
Appendix E Detailed Experiment Results
We present detailed comparisons between RL models and baselines.
@@ Appendix E Detailed Experiment Results
Dual BD 3 bins 0.000 71.4 1.734 0.123 55.2 57.1
Dual BD 5 bins 0.000 96.3 2.383 0.265 39.7 26.4
Dual BD 10 bins 0.000 92.3 4.673 0.309 42.4 13.3
@@
Dual BD 3 bins 0.001 82.2 1.614 0.110 46.9 55.5
Dual BD 5 bins 0.001 93.6 2.281 0.217 41.3 27.2
Dual BD 10 bins 0.001 96.6 4.445 0.292 39.1 14.1
@@
Dual BD 3 bins 0.010 76.8 1.576 0.127 53.1 56.3
Dual BD 5 bins 0.010 84.9 2.091 0.119 43.9 40.3
Dual BD 10 bins 0.010 92.8 4.072 0.245 41.4 16.6
@@
Appendix F Model Implementation and Training
F.1 Double Q-learning
@@
F.2 Binary vp1 Model
The Binary $vp_{1}$ model represents the simplest action space, treating vasopressin administration as a binary decision (on/off). The $vp_{2}$ levels are used as part of the state. Training employs standard Q-learning TD loss with conservative logsumexp regularization over both action choices to prevent overestimation of out-of-distribution actions. The model processes states through fully connected networks with ReLU activations, using double Q-learning with target networks for stability.
F.3 Dual Mixed Model
@@
F.4 Dual Block Discrete Model
@@
F.5 Dual Stepwise Model
@@
F.6 LSTM Block Discrete Model
@@
Appendix G Model Architectures
The model architectures for general Q-learning networks is shown in Table 7. The model architectures for LSTM Block Discrete networks is shown in Table 8.
@@ Appendix G Model Architectures
Appendix H Hyper-parameter Configurations
The shared hyper-parameters are shown in Table 9. The hyper-parameters specific to different action space models are shown in Table 10.
@@ Appendix H Hyper-parameter Configurations
Epochs 100 Number of training epochs
Validation Batches 10 Batches used for validation
Random Seed 42 Ensures reproducibility